OpenStudy (anonymous):

A circular pool, 24 1/2 foot in diameter, is installed onto a rectangular yard (42 by 28 feet). The yard needs new sod. How much sod should be ordered to cover the yard? (use 3.14 and round 2 decimal places)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we have to compute the area of the yard, namely: A = 42*28=...?

OpenStudy (anonymous):

A=1176

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, we have to compute the area A of the circular pool, namely: \[\large A = 3.14 \times {R^2}\] where, as usual, R is the radius, and R= D/2, D is the diameter

OpenStudy (anonymous):

A=3.14* 1176^2?

OpenStudy (michele_laino):

the diameter is D= 49/2 = 24.5 feet so the radius is: R= 24.5/2= 12.25 finally A= 3.14*12.25^2=...?

OpenStudy (anonymous):

471.19625

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

finally, your answer is given by the subsequent difference: \[\large 1176 - 471.19 = ...?\]

OpenStudy (anonymous):

Nice! you are so smart!

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

yes!:)

OpenStudy (michele_laino):

:)